# Enhancing Ultra High Resolution Remote Sensing Imagery Analysis with ImageRAG

### v2

@@ VI-D Result of Visual Cue Retrieval Task
From Table IV, we can see that introducing the slow path mechanism boosts the mean recall for both T=0.1 and T=0.3, with even higher mean recall at T=0.1 when both LRSD and CRSD are used.
However, the result is not significant because we force the model to output visual cues even if they do not meet the $\epsilon$ threshold, which hurt the result under this metric. However, if we did not use this metric, the comparison of the mean recall will not be fair. The reason is, in practice, we find that both the vanilla and inferring models are sensitive to the accuracy of the retrieved visual cues. Inaccurate visual cues can sometimes lead to worse ImageRAG performance than the baseline. In Table III, we demonstrate this phenomenon: a noticeable (-26.67%) decline is happened when we replace the ground truth ROI box with a random box. Therefore, during deployment, we adopt a conservative strategy: we do not provide the visual cue if the retrieved result is not believable (providing incorrect information is worse than providing none). That is, we filter out visual cues with lower confidence, which enhances the overall framework’s accuracy.
When the Slow path is disabled, the Fast path is used less frequently (10.67% utilization). The model tends to answer questions without ImageRAG, in a zero-shot manner. As the Fast path is uncertain, the framework often outputs void visual cues. However, when the Slow path is applied, the model uses ImageRAG more frequently to answer questions (45.33% utilization).
@@ VII Ablation Study
In this ablation study, we examine the key factors that influence ImageRAG’s performance, such as the Patch Division Algorithm, the Image and Text Encoder, and the Proxy Selection Function. In these studies, we use an ensemble of results choose from different thresholds to study: $\epsilon\in\{0.3,0.5,0.7\}$, $\delta_{1}\in\{0.1,0.2,0.3\}$, and $\delta_{2}\in\{0.3,0.5,0.7\}$. The rationale behind this approach is we aim to limit the return of inaccurate visual cues, particularly for $\delta_{1}$, as it controls the output from querying LRSD database.
Additionally, we explore how the performance of the inferring model changes when the ROI size is enlarged, which may cause the evidence to become more vague.
@@ VII-B Patch Division Algorithm
TABLE V: Regular VQA Task result with different patch division algorithms.
@@
Table V presents the performance of different patch division algorithms. "Complete Cover" achieves the highest average Counting and Color score, primarily attributed to its diverse patch scales. "Cascade Grid" offers the best average Position accuracy, likely due to its effective balance between patch scales and the number of distractor patches (better than dense "Complete Cover"). ViT demonstrates the most balanced performance, likely due to the small number of patches reducing potential distractions.
@@ VII-C Proxy Selection Function
TABLE VI: Regular VQA Task Result with different proxy selection functions.
@@
Table VI shows the performance of different proxy feature selection functions. Overall, their performance is comparable, yet each excels in specific tasks: "Reranking" in Position, "Clustering" in Color, and "Prototype" in Counting. In practice, high-accuracy trails mostly come from "Prototype" and "Clustering". However, since "Clustering" is much more time-consuming, "Prototype" is more recommended.
@@ VIII Result Visualization and Discussion
IX Related Work
IX-A Remote Sensing Vision-Language Models
@@
Remote Sensing Vision-Language Models (RSVLMs) are developed to process and analyze geospatial data by incorporating both visual and linguistic information [59]. These models are pre-trained on large-scale remote sensing imagery and relevant text, enabling them to adapt to various VLM tasks in RS domains. RSVLMS have demonstrated promising results in specialized applications, including scene classification [4, 9], object detection [3, 5], semantic segmentation [60, 61], image captioning [4, 3], text-image retrieval [8, 9], visual grounding [3, 50, 5], image generation [8, 62, 63], etc. Many of them are capable to complete multiple downstream tasks.
@@
RSVLMs can be categorized into three types based on their input-output mechanisms  [64]. Contrastive RSVLMs process both text and images as inputs, generating similarity scores essential for tasks such as image-text retrieval and zero-shot scene classification. Conversational RSVLMs also take text and images as inputs but produce textual responses, leveraging LLMs for tasks like captioning and visual question answering. In contrast, generative RSVLMs are conditioned on either text or images to generate synthetic remote sensing images, typically employing conditional diffusion processes for controlled image synthesis.
@@
For contrastive VLGFMs, research primarily focuses on extending CLIP [25] to RS applications, emphasizing the development of RS-specific datasets and benchmarks. Liu et al. proposed RemoteCLIP [9], a vision-language foundation model for remote sensing that achieves significant improvements in various downstream tasks through multi-task pre-training and data expansion. Zhang et al. developed GeoRSCLIP [8] based on the large-scale remote sensing image-text dataset RS5M, which demonstrates excellent performance in zero-shot classification, cross-modal text-image retrieval, and semantic localization tasks. Mall et al. introduced GRAFT [65], a method to train vision-language models for satellite images without textual annotations by using ground images as an intermediary. Wang et al. constructed SkyScript [66], a large and semantically diverse vision-language dataset for remote sensing, and developed SkyCLIP through continual pre-training, which shows superior performance in zero-shot scene classification, fine-grained attribute classification, and cross-modal retrieval tasks.
@@
IX-B Vision-Language Model with Visual Cues
@@
IX-C Retrieval-Augmented Generation
@@
The effectiveness of RAG systems heavily depends on the quality and relevance of the retrieved knowledge, which directly influences the accuracy and factual grounding of generated content. To enhance retrieval efficiency and overcome the limitations of traditional methods, several advancements have been proposed, particularly for zero-shot and few-shot retrieval tasks. Techniques such as HyDE [89] and REINA [90] utilize LLMs to generate hypothetical documents, improving retrieval performance without requiring labeled data. The Rewrite-Retrieve-Read  [91] framework introduces a query rewriting step, allowing the input query to be better aligned with retrieval modules. By using reinforcement learning to adapt queries, R3 enhances retrieval quality, improving performance in open-domain and multiple-choice question answering tasks. Promptagator [92] demonstrates the effectiveness of few-shot learning in dense retrieval, utilizing LLMs to generate synthetic training data from minimal examples, surpassing models trained on large-scale datasets like MS MARCO. This underscores the viability of few-shot learning and LLM-generated synthetic data in resource-constrained settings. To bridge the preference gap between retrievers and LLMs, Zixuan Ke et al. [93] introduce the BGM framework, which employs a sequence-to-sequence model to align retrieved information with LLM preferences. BGM optimizes the retriever-LLM interaction, enhancing performance in downstream tasks such as question answering and personalized text generation.
@@
Methods like RECITE [94] and ITER-RETGEN [95] focus on improving factual accuracy and integrating retrieved knowledge, ensuring grounded, accurate responses through knowledge recitation and iterative retrieval-generation. Frameworks such as Selfmem [96] and Self-RAG [97] enhance factual consistency via self-reflection. Techniques like Step-Back Prompting [98] improve reasoning by guiding LLMs to abstract high-level concepts. The GENREAD approach [99] replaces traditional retrieval with LLM-generated contextual documents, demonstrating superior performance in knowledge-intensive tasks like open-domain QA and fact-checking.
@@
Iterative and active retrieval-generation strategies aim to dynamically enhance both retrieval and generation processes. ITER-RETGEN [95] alternates between retrieval and generation to improve response quality, while FLARE [100] predicts and adapts to future information needs during generation. These methods boost relevance and accuracy in knowledge-intensive tasks. The Adaptive-RAG framework[101] selects retrieval strategies based on query complexity, outperforming static models in efficiency and accuracy.
@@
Reducing hallucinations, where models generate factually incorrect or fabricated information, is a critical challenge in RAG systems. To address this, various strategies focus on improving the trustworthiness of generated outputs by leveraging more reliable retrieval mechanisms and robust post-generation processes. Methods such as RAGTruth [102] introduce large-scale datasets designed to detect and mitigate hallucinations, enabling models to generate more trustworthy responses. Ayala et al.  [103] use external knowledge sources to reduce errors in structured output generation, enhancing the reliability of RAG systems in practical applications.
@@ IX-D Multimodal RAG
The latest advancements in RAG in the image domain have led to significant improvements. RA-CM3 [108] enhances both text-to-image and image-to-text generation by combining the CLIP [109] retriever and the CM3 Transformer generator, achieving a performance boost while reducing computational costs by over 30%. Mortaheb et al. [107] introduced a re-ranking mechanism based on a relevance score model [110], improving context selection during retrieval and reducing hallucinations, thereby enhancing the quality of generated responses. Yu et al.’s VisRAG [111] framework bypasses the text parsing stage to directly process multi-modal documents containing both text and images, achieving substantial improvements in multi-modal tasks. Bonomo and Bianco’s Visual RAG [112] expands the visual knowledge of large MLLMs without the need for fine-tuning by dynamically retrieving relevant examples, offering high computational efficiency. Finally, Riedler and Langer’s work on multimodal inputs for industrial applications demonstrates that integrating both images and text in RAG systems significantly improves performance [113].
RAG has also made significant progress in the video domain, driving innovations in long video comprehension. Yongdong Luo et al. propose Video-RAG [114], which enhances long video understanding by integrating visually-aligned auxiliary texts into large video-language models, surpassing models like Gemini1.5-Pro and GPT-4o. Jeong et al. introduce VideoRAG[115], a method that dynamically retrieves relevant videos based on user queries and combines both visual and textual information to generate more contextually rich responses, showing marked improvements over traditional RAG approaches. Ma et al. present DrVideo[116], a system that converts long videos into text-based documents and iteratively retrieves missing information, achieving high accuracy in key frame identification. Arefeen et al. introduce iRAG[117], an incremental method that accelerates video-to-text processing by using lightweight models for fast indexing and heavyweight models for detailed extraction, making it well-suited for real-time video analysis. Lastly, Zhang et al. propose OmAgent[118], a multi-modal agent framework that minimizes information loss in video understanding tasks by dynamically invoking retrieval tools, thereby enhancing reasoning and event localization.
X Conclusion and Future Work

### v3

@@ VI-D Result of Visual Cue Retrieval Task
From Table IV, we can see that introducing the slow path mechanism boosts the mean recall for both T=0.1 and T=0.3, with even higher mean recall at T=0.1 when both LRSD and CRSD are used. However, the result is not significant because we force the model to output visual cues even if they do not meet the $\epsilon$ threshold, which hurt the result under this metric. However, if we did not use this metric, the comparison of the mean recall will not be fair. The reason is, in practice, we find that both the vanilla and inferring models are sensitive to the accuracy of the retrieved visual cues. Inaccurate visual cues can sometimes lead to worse ImageRAG performance than the baseline. In Table III, we demonstrate this phenomenon: a noticeable (-26.67%) decline is happened when we replace the ground truth ROI box with a random box. Therefore, during deployment, we adopt a conservative strategy: we do not provide the visual cue if the retrieved result is not believable (providing incorrect information is worse than providing none). That is, we filter out visual cues with lower confidence, which enhances the overall framework’s accuracy.
When slow path is disabled, fast path is used less frequently (10.67% utilization). The model tends to answer questions without ImageRAG, in a zero-shot manner. As fast path is uncertain, the framework often outputs void visual cues. However, when slow path is applied, the model uses ImageRAG more frequently to answer questions (45.33% utilization).
VI-E Result Robustness on Larger Dataset
In Section VI-B, ImageRAG showed remarkable performance on the MME-RealWorld-Lite-RS dataset. However, this dataset is relatively small, containing only 150 samples in total, with 50 samples per subtask. To test how ImageRAG performs with larger datasets, we created a series of datasets with different volumes.
We add 360 more samples (120 per subtask) from MME-RealWorld-RS, following the labeling standards outlined in section II-B. These new samples do not overlap with those in MME-RealWorld-Lite-RS and were examined by human to make sure the question and answer are not ambiguous nor vague. We refer to the original MME-RealWorld-Lite-RS dataset as $\mathbb{D}_{150}$. We also present $\mathbb{D}_{300}$, $\mathbb{D}_{450}$, and $\mathbb{D}_{510}$, each with an equal number of samples for each subtask. Moreover, each smaller dataset is a subset of the larger ones. This means $\mathbb{D}_{150}\subseteq\mathbb{D}_{300}\subseteq\mathbb{D}_{450}\subseteq\mathbb{D}_{510}$.
TABLE V: Performance of ImageRAG on larger datasets.
Dataset Accuracy
Position Color Count Average
$\mathbb{D}_{150}$ 64.00 62.00 30.00 52.00
$\mathbb{D}_{300}$ 54.00 65.00 33.00 50.67
$\mathbb{D}_{450}$ 58.67 63.33 32.00 51.33
$\mathbb{D}_{510}$ 59.41 64.12 32.35 51.96
As shown in Table V, ImageRAG demonstrates consistent performance across datasets of varying sizes, indicating that it is robust on larger datasets.
@@ VII Ablation Study
In this ablation study, we examine the key factors that influence ImageRAG’s performance, such as the Patch Division Algorithm, the Image and Text Encoder, and the Proxy Selection Function. We use an ensemble of results choose from different thresholds to study: $\epsilon\in\{0.3,0.5,0.7\}$, $\delta_{1}\in\{0.1,0.2,0.3\}$, and $\delta_{2}\in\{0.3,0.5,0.7\}$. The rationale behind this setting is we aim to limit the return of inaccurate visual cues, particularly for $\delta_{1}$, as it controls the output from querying LRSD database. Additionally, we explore how the performance of the inferring model changes when the ROI size is enlarged, which may cause the evidence to become more vague.
@@ VII-B Patch Division Algorithm
TABLE VI: Regular VQA Task result with different patch division algorithms.
@@
Table VI presents the performance of different patch division algorithms. "Complete Cover" achieves the highest average Counting and Color score, primarily attributed to its diverse patch scales. "Cascade Grid" offers the best average Position accuracy, likely due to its effective balance between patch scales and the number of distractor patches (better than dense "Complete Cover"). ViT demonstrates the most balanced performance, likely due to the small number of patches reducing potential distractions.
@@ VII-C Proxy Selection Function
TABLE VII: Regular VQA Task Result with different proxy selection functions.
@@
Table VII shows the performance of different proxy feature selection functions. Overall, their performance is comparable, yet each excels in specific tasks: "Reranking" in Position, "Clustering" in Color, and "Prototype" in Counting. In practice, high-accuracy trails mostly come from "Prototype" and "Clustering". However, since "Clustering" is much more time-consuming, "Prototype" is more recommended.
@@ VIII Result Visualization and Discussion
IX Computational Resource
The LoRA fine-tuning process took 2 hours for 2 epochs using 8 NVIDIA A100-80GB GPUs  [^9] https://internvl.readthedocs.io/en/latest/internvl2.5/finetune.html, and hosting Qwen-32B-Instruct required 4 NVIDIA A100-40GB GPUs. As for running ImageRAG, a single NVIDIA A100-40GB or NVIDIA RTX 4090-24GB would be enough.
X A Cookbook for Adapting ImageRAG to Different Image Modalities and Domains
In this paper, we focus on the optical ultra high resolution remote sensing image, but ImageRAG is a geneal framework that is able to be applied in different UHR image modalities (e.g. SAR and Hyperspectral data), and image domains (e.g. Medical Imaging). Several modules can be redesigned to better fit the specific task, and some others can remain invariant.
X-A Patchify
In section V-A, we introduced three selected patch division algorithm: "ViT", "Complete Cover", "Cascade Grid". These are model-agnostic approaches, which means it can be adopted to different image modalities and domains without any modification or retraining.
X-B Question Analyzing Module
We deploy Qwen2.5-32B on SGLang for API serving to extract the target-of-interest from questions, ensuring the extracted key phrases are as accurate as possible. However, since the target-of-interest extraction task is relatively simple, this model could potentially be replaced by lighter general LLMs such as MiniCPM3-4B [64] and Qwen3-0.6B in real-world deployment scenarios, or domain-specific LLM like Med-PaLM 2 [65] when dealing with UHR medical images.
X-C Text-Image Retrieval Module
This module is of critical importance to the ImageRAG framework, since image patches, target-of-interest text, image data from vector database are all encoded as feature vectors using the VLM within this module. More importantly, the visual cues are retrieved based on the similarities between these features. In our study, we tried domain-specific contrastive VLMs (RemoteCLIP, GeoRSCLIP) and general VLMs (CLIP, MCIPCLIP) for UHR optical remote sensing images. For encoding SAR, Multispectral, infrared or hyperspectral imaging alongside corresponding text, GeoLangBind [66] can be considered. As for medical imaging from other image domain, MedCLIP [67] can be chosen.
X-D Text-Text Retrieval Module and Vector Database
This module matches the data from the vector database with the key phrases extracted from the question. Sentence-BERT is generally suitable for text-text retrieval. When dealing with modality-specific or domain-specific data, it is advisable to collect images accompanied by text labels or captions to construct the vector database effectively. In scenarios where access the image-text paired data is limited, the calculation of proxy image embeddings can be replaced by leveraging all available data through in-context few-shot learning [68], which involves constructing task examples alongside the question to send to the VLM. This approach aligns with the ultimate goal of the slow path, which is to select useful image patches with the aid of external data as visual cues to send to the MLLM.
XI Limitation
The ImageRAG framework has two major limitations. First, the time cost is significant. Second, the performance heavily depends on the accuracy of retrieved visual cues.
TABLE VIII: We compared the time cost of ImageRAG across different strategies. "Mode" shows whether ImageRAG used only the fast path or a combination of fast and slow paths as detailed in Section IV-A. "Vector Database" reflects if LRSD and CRSD were used as described in Section V-D. "Time Cost" represents the average processing time per question in seconds. All patches and their corresponding features are pre-extracted and used as caches. Vector Databases are pre-built.
Mode Vector Database Time Cost Note
Baseline - 1.14 sec Direct Inference
Fast Path Only - 2.70 sec All Fast Path
Fast + Slow Path LRSD 2.89 sec All Slow Path
Fast + Slow Path LRSD + CRSD 3.02 sec All Slow Path
Fast + Slow Path LRSD + CRSD 2.81 sec Fast & Slow Path
As shown in Table VIII, we compared the relative time cost of ImageRAG across different strategies. For fair comparison, we pre-extracted all patches and their corresponding features and used them as caches (This process is parallelized using Ray  [^10] https://github.com/ray-project/ray with multiple GPUs). Vector Databases are pre-built. Compared with the baseline setting, which does not use ImageRAG, the additional time cost primarily comes from question analyzing and visual cue retrieval. When using the fast path for all questions, the time cost doubles (1.14 seconds compared to 2.70 seconds). If all questions require the worst-case scenario—where the fast path fails and the slow path must search both vector databases to find visual cues—the time cost is almost tripled (1.14 seconds compared to 3.02 seconds). In practical scenarios, when mixing fast and slow paths across the dataset, the time cost is nearly 2.5 times higher than the baseline (1.14 seconds compared to 2.81 seconds). However, the absolute time cost should be acceptable in practical development as the analysis processing time remains measured in seconds. To further enhance speed, the Qwen2.5-32B model used for question processing can be replaced with a lighter model as mentioned in section X-B.
Moreover, as shown in Section VI-C and Section VII-D, the performance of ImageRAG is significantly affected by the accuracy and size of the retrieved visual cues. When the retrieved visual cues are overly large or incorrect, the performance of ImageRAG can greatly decline (Figure 9) and may even drop below the baseline (as seen in the "Random Box" row of Table III). Techniques referenced in section XII-C that improved the accuracy of the retrieved results of RAG can also be applied to mitigate this issue.
XII Related Work
XII-A Remote Sensing meets Vision-Language Models
@@
Remote Sensing Vision-Language Models (RSVLMs) are developed to analyze geospatial data by incorporating both visual and linguistic information [69]. These models are pre-trained on large-scale RSI and relevant text, enabling them to adapt to various VLM tasks in RS domains. RSVLMS have demonstrated promising results in specialized applications, including scene classification [4, 9], object detection [3, 5], semantic segmentation [70, 71], image captioning [4, 3], text-image retrieval [8, 9], visual grounding [3, 52, 5], image generation [8, 72, 73], etc. Many of them are capable to complete multiple downstream tasks.
@@
RSVLMs can be categorized into three types based on their input-output mechanisms  [74]. Contrastive RSVLMs process both text and images as inputs, generating similarity scores essential for tasks such as image-text retrieval and zero-shot scene classification. Conversational (generative) RSVLMs also take text and images as inputs but produce textual responses, leveraging LLMs for tasks like captioning and visual question answering. Besides, some RSVLMs are conditioned on either text or images to generate synthetic remote sensing images, typically employing conditional diffusion processes for controlled image synthesis.
@@
For contrastive RSVLMs, research primarily focuses on extending CLIP [27] to RS applications, emphasizing the development of RS-specific datasets and benchmarks. Liu et al. proposed RemoteCLIP [9], a vision-language foundation model for remote sensing that achieves significant improvements in various downstream tasks through multi-task pre-training and data expansion. Zhang et al. developed GeoRSCLIP [8] based on the large-scale remote sensing image-text dataset RS5M, which demonstrates excellent performance in zero-shot classification, cross-modal text-image retrieval, and semantic localization tasks. Mall et al. introduced GRAFT [75], a method to train vision-language models for satellite images without textual annotations by using ground images as an intermediary. Wang et al. constructed SkyScript [76], a large and semantically diverse vision-language dataset for remote sensing, and developed SkyCLIP through continual pre-training, which shows superior performance in zero-shot scene classification, fine-grained attribute classification, and cross-modal retrieval tasks.
@@
XII-B Vision-Language Model with Visual Cues
@@
XII-C Retrieval-Augmented Generation
@@
The effectiveness of RAG systems heavily depends on the quality and relevance of the retrieved knowledge, which directly influences the accuracy and factual grounding of generated content. To enhance retrieval efficiency and overcome the limitations of traditional methods, several advancements have been proposed, particularly for zero-shot and few-shot retrieval tasks. Techniques such as HyDE [99] and REINA [100] utilize LLMs to generate hypothetical documents, improving retrieval performance without requiring labeled data. The Rewrite-Retrieve-Read  [101] framework introduces a query rewriting step, allowing the input query to be better aligned with retrieval modules. By using reinforcement learning to adapt queries, R3 enhances retrieval quality, improving performance in open-domain and multiple-choice question answering tasks. Promptagator [102] demonstrates the effectiveness of few-shot learning in dense retrieval, utilizing LLMs to generate synthetic training data from minimal examples, surpassing models trained on large-scale datasets like MS MARCO. This underscores the viability of few-shot learning and LLM-generated synthetic data in resource-constrained settings. To bridge the preference gap between retrievers and LLMs, Zixuan Ke et al. [103] introduce the BGM framework, which employs a sequence-to-sequence model to align retrieved information with LLM preferences.
@@
Methods like RECITE [104] and ITER-RETGEN [105] focus on improving factual accuracy and integrating retrieved knowledge, ensuring grounded, accurate responses through knowledge recitation and iterative retrieval-generation. Frameworks such as Selfmem [106] and Self-RAG [107] enhance factual consistency via self-reflection. Techniques like Step-Back Prompting [108] improve reasoning by guiding LLMs to abstract concepts. The GENREAD approach [109] replaces traditional retrieval with LLM-generated contextual documents, demonstrating superior performance in knowledge-intensive tasks like open-domain QA and fact-checking.
@@
Iterative and active retrieval-generation strategies aim to dynamically enhance both retrieval and generation processes. ITER-RETGEN [105] alternates between retrieval and generation to improve response quality, while FLARE [110] predicts and adapts to future information needs during generation. These methods boost relevance and accuracy in knowledge-intensive tasks. The Adaptive-RAG framework[111] selects retrieval strategies based on query complexity, outperforming static models in efficiency and accuracy. FunnelRAG [112] proposes a progressive retrieval paradigm with coarse-to-fine granularity for RAG, to enable load balancing and improve retrieval performance.
@@
Reducing hallucinations in response is a critical challenge in RAG systems. To address this, various strategies focus on improving the trustworthiness of generated outputs by leveraging more reliable retrieval mechanisms and robust post-generation processes. Methods such as RAGTruth [113] introduce large-scale datasets designed to detect and mitigate hallucinations, enabling models to generate more trustworthy responses. SEER [114] proposes a novel evidence extraction learning paradigm, which utilizes the model to calibrate its extraction preference via self-alignment. Ayala et al.  [115] use external knowledge sources to reduce errors in structured output generation, enhancing the reliability of RAG systems in practical applications.
@@ XII-D Multimodal RAG
The latest advancements in RAG in the image domain have led to significant improvements  [120]. RA-CM3 [121] enhances both text-to-image and image-to-text generation by combining the CLIP retriever and the CM3 Transformer generator, achieving a performance boost while reducing computational costs by over 30%. Mortaheb et al. introduced a re-ranking mechanism based on a relevance score model [122], improving context selection during retrieval and reducing hallucinations, thereby enhancing the quality of generated responses. Yu et al.’s VisRAG [123] framework bypasses the text parsing stage to directly process multi-modal documents containing both text and images, achieving substantial improvements in multi-modal tasks. Bonomo and Bianco’s Visual RAG [124] expands the visual knowledge of large MLLMs without the need for fine-tuning by dynamically retrieving relevant examples, offering high computational efficiency. Riedler and Langer’s work on multimodal inputs for industrial applications demonstrates that integrating both images and text in RAG systems significantly improves performance [125].
RAG has also made significant progress in the video domain, driving innovations in long video comprehension. Yongdong Luo et al. propose Video-RAG [126], which enhances long video understanding by integrating visually-aligned auxiliary texts into large video-language models, surpassing models like Gemini1.5-Pro and GPT-4o. Jeong et al. introduce VideoRAG[127], a method that dynamically retrieves relevant videos based on user queries and combines both visual and textual information to generate more contextually rich responses, showing marked improvements over traditional RAG approaches. Ma et al. present DrVideo[128], a system that converts long videos into text-based documents and iteratively retrieves missing information, achieving high accuracy in key frame identification. Arefeen et al. introduce iRAG[129], an incremental method that accelerates video-to-text processing by using lightweight models for fast indexing and heavyweight models for detailed extraction, making it well-suited for real-time video analysis. Zhang et al. propose OmAgent[130], a multi-modal agent framework that minimizes information loss in video understanding tasks by dynamically invoking retrieval tools, thereby enhancing reasoning and event localization.
XIII Conclusion and Future Work